OpenStudy (meilendurcer):

Find the length of the unknown side. Round your answer to the nearest whole number. A. 7 B. 9 C. 10 D. 12 2. The numbers 20,21,29. form a Pythagorean triple. Which other set of numbers is a Pythagorean triple? A. 21,22,30 B. 40,42,58 C. 60,66,87 D. 70,84,116

OpenStudy (anonymous):

what is the figure? for number one?

OpenStudy (mathmate):

For #2, all multiples of Pythagorean triples remain Pythagorean triples, for example 3,4,5 and 6,8,10 are Pythagorean triples.

OpenStudy (anonymous):

the missing side is the hypotenuse so your equation would be \[a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (anonymous):

or called at pathagory thereom

OpenStudy (anonymous):

so it would be \[a ^{2}+b ^{2}=c ^{2} ...7^{2}+7^{2}=c ^{2}...c ^{2}= 14^{2}\]

OpenStudy (anonymous):

now 14 times 14 = 196 and since its squared you need to find the square root

OpenStudy (mathmate):

Ouch! \( 7^2 + 7^2 = 49+49 =98 \ne 14^2 \)